OpenStudy (anonymous):

a particle is projected with velocity 9m/s up a line of greatest slope of a smooth plane inclined at an angle a to the horizontal where sin a=4/7. find the speed of the particle when it has travelled 5m up the plane.

OpenStudy (y2o2):

get the a where a = -gsinθ = 9.81 * (4/7) = -5.605 m/s² so use v² = u² + 2as

OpenStudy (anonymous):

i get the answer 4.99m/s.is it correct?

OpenStudy (dumbcow):

yes that is correct